a car travels between a and b at an average speed of 60km/h. if the car increases its average speed to 100 km/h would take 10 minutes less to make the trip. How far it is between the towns?
Answer by mananth(16946)

Time difference with increased speed = 10 minutes = 10/60 = 1/6 hours


let distance be x
x/ 60 -x/100 = 1/6
multiply equation by 600
10x-6x=100
4x=100
/4
x=25 km